OpenStudy (anonymous):

After the fraction x+1/3 - x+2/2x have been combined using the Least Common Denominator of 6x, what is the numerator?

OpenStudy (anonymous):

could this be \[\large \frac{x+1}{3}-\frac{x+2}{2x}\]?

OpenStudy (anonymous):

Do I first multiply the first fraction by 2x/2x and the second fraction by 3/3? and yes that is the expression :)

OpenStudy (anonymous):

yes that is what you do exactly

OpenStudy (anonymous):

I get 2x^2+1/6x - 3x+2/6x?

OpenStudy (anonymous):

hmm no i don't think so

OpenStudy (anonymous):

Should I simplify to 2x^2+3x+3/6x

OpenStudy (anonymous):

first of all you have too many fraction bars there, i assume you meant \[\frac{2x^2+1}{6x}-\frac{3x+2}{6x}\] but that is also wrong, you have to distribute when you multiply

OpenStudy (anonymous):

Would you show me? Please :)

OpenStudy (anonymous):

\[\large \frac{x+1}{3}-\frac{x+2}{2x}\] \[\frac{2x}{2x}\times \frac{x+1}{3}-\frac{3}{3}\frac{x+2}{2x}\] is right but you have to think of it as \[\frac{2x}{2x}\times \frac{(x+1)}{3}-\frac{3}{3}\times \frac{(x+2)}{2x}\]

OpenStudy (anonymous):

Ah, so I should get 2x^2 + 2x for the first fractions numerator, 6x for the first fractions denominator?

OpenStudy (anonymous):

the numerator and and denominator of a fraction carry their own parentheses even if you don't see them think "order of operations" when you multiply you will get \[\frac{2x^2+2x-3x-6}{6x}\]

OpenStudy (anonymous):

Then simplify right? To 2x^2 -1x - 6 over 6x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or keep it, no simplification

OpenStudy (anonymous):

okay thanks!! :):)

OpenStudy (anonymous):

i would call that "combine like terms" there is no such mathematical operation as "simplify" no matter how many times your math teacher says it it is not real

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Haha, I understand and thank you really, much appreciated! Very helpful c:::

OpenStudy (anonymous):

thanks